12 Electric field and forces at various locations Hidden source charges produce the illustrated electric field. a) Determine the force exerted on a proton that is placed at 8 x 103 N/C 4x 103 N/C A. Determine the acceleration of a proton at this point. → b) Determine the force exerted on an electron that is placed at B. Determine the acceleration of an electron at this point. c) Determine the force exerted on an electron that is placed at C. Determine the acceleration of an electron at this point. _ ..> -→
Answer #1

Magnitude of force on charge q, in electric fueld E

= Eq   

Direction of force on +ve charge is in the direction of electric field and that on -ve charge is in the direction opposite to direction of electric field.

Charge of proton =+ 1.6×10^-19 C

Charge of electron = - 1.6×10^-19 C

Electric field at B and A = 8×10^3 N/C

Electric field at C = 4×10^3 N/C

a) Force on proton at A= Ea * e = 8×10^3(N/C)*1.6×10^-19(C)

= 1.3×10^-15 N in the direction of field at A

Acceleration = force / mass

Acceleration of proton at A = 1.3×10^-15(N)/1.67×10^-27(kg)

= 7.7×10^11 m/s^2

b) Force on electron at B = Eb*e = 8×10^3(N/C)*1.6×10^-19(C)

=1.3×10^-15 N opposite to direction of field at B

Acceleration of electron at B = force on electron at B / mass of electron = 1.3×10^-15(N/C)/9.1×10^-31(kg) = 1.4×10^15 m/s^2

c) Force on electron at C = Ec*e = 6.4×10-16 N opposite to direction of field at C

Acceleration of electron at C = force at C/ mass of electron

= 7.0×10^14 m/s^2